On February 1, 2017, Marsh Contractors agreed to construct a building at a contract price of $5,720,000. Marsh estimated total construction costs would be $3,936,000 and the project would be finished in 2019. Information relating to the costs and billings for this contract is as follows:

2017 2018 2019
Total costs incurred to date $1,476,000 $2,592,000 $4,560,000
Estimated costs to complete 2,460,000 1,728,000 -0-
Customer billings to date 2,120,000 3,936,000 5,520,000
Collections to date 1,920,000 3,420,000 5,420,000


Fill in the correct amounts on the following schedule. For percentage-of-completion accounting and for completed-contract accounting, show the gross profit that should be recorded for 2017, 2018, and 2019.

Percentage-of-Completion
Gross Profit
Completed-Contract
Gross Profit
2017 $

2017 $

2018 $

2018 $

2019 $

2019 $
Answer #1

Answer

  • Working for Percentage of Completion method

Working

Year end 2017

Year end 2018

Year end 2019

A

Contract Price

$       5,720,000.00

$       5,720,000.00

$          5,720,000.00

B

Cost Incurred to Date

$       1,476,000.00

$       2,592,000.00

$          4,560,000.00

C

Estimated cost yet to be incurred to complete the contract

$       2,460,000.00

$       1,728,000.00

$                                -  

D = B+C

Total Cost

$       3,936,000.00

$       4,320,000.00

$          4,560,000.00

E = (B/D) x 100

% of Completion

37.50%

60.00%

100.00%

F = A x E

Revenue to date

$       2,145,000.00

$       3,432,000.00

$          5,720,000.00

G

Revenue of Previous year

$                            -  

$       2,145,000.00

$          3,432,000.00

H = F - G

Net Revenue this year

$       2,145,000.00

$       1,287,000.00

$          2,288,000.00

I (=B)

Cost to date

$       1,476,000.00

$       2,592,000.00

$          4,560,000.00

J

Cost to date of previous year

$                            -  

$       1,476,000.00

$          2,592,000.00

K = I - J

Net Cost for the year

$       1,476,000.00

$       1,116,000.00

$          1,968,000.00

L = H - K

Gross Profits

$           669,000.00

$          171,000.00

$              320,000.00

  • Requirements

Percentage-of-Completion
Gross Profit

Completed-Contract
Gross Profit

2017

$ 669000

phpMLlOcS.png

2017

$0

phpeejAyx.png

2018

$ 171000

phpoXPG8X.png

2018

$ 0

phpIj8UJw.png

2019

$ 320000

phpwQ6BrM.png

2019

$ 1160000 [$5720000 – 4560000 cost incurred]

php6VnxBs.png